# Pathway to Secure and Trustworthy 6G for LLMs: Attacks, Defense, and Opportunities

### v1

@@ Abstract
Recently, large language models (LLMs) have been gaining a lot of interest due to their adaptability and extensibility in emerging applications, including communication networks. It is anticipated that 6G mobile edge computing networks will be able to support LLMs as a service, as they provide ultra reliable low-latency communications and closed loop massive connectivity. However, LLMs are vulnerable to data and model privacy issues that affect the trustworthiness of LLMs to be deployed for user-based services. In this paper, we explore the security vulnerabilities associated with fine-tuning LLMs in 6G networks, in particular the membership inference attack. We define the characteristics of an attack network that can perform a membership inference attack if the attacker has access to the fine-tuned model for the downstream task. We show that the membership inference attacks are effective for any downstream task, which can lead to a personal data breach when using LLM as a service. The experimental results show that the attack success rate of maximum 92% can be achieved on named entity recognition task. Based on the experimental analysis, we discuss possible defense mechanisms and present possible research directions to make the LLMs more trustworthy in the context of 6G networks.
@@ I Introduction
The emergence of attention networks has been a stepping stone for transformer architectures, which also led to the introduction of large language models (LLMs). More recently, LLMs are seen as the most significant advance in the field of artificial intelligence (AI) and a potential pathway to artificial general intelligence (AGI) [1]. Every tech giant is in a race to advance in the field of LLMs by leveraging generative AI (GAI). Notable examples of LLMs from the tech giants are GPT-4 from OpenAI, LLaMA-3 from Meta and PALM from Google. However, there are also new players in this field that surpass the performance of the LLMs mentioned above. These include Mistral (in collaboration with NVIDIA), DCLM from Apple, xLAM from Salesforce, v2 chat from Deepseek, Groq, Claude, SmolLM and many more. These LLMs are trained on diverse and large amounts of datasets scraped or curated from the Internet. Some LLMs focus on increasing model size, such as GPT, while others find new ways to improve the generalization of LLMs through data curation, model quantization, and innovative techniques. Examples of such LLMs are Claude, LLaMA, DCLM and Groq, which have recently outperformed GPT on various language tasks. In continuation of the above-mentioned advances in LLMs, several enterprises are leveraging pre-trained encoders of LLMs to varying degrees to develop their own customized solutions for various applications and sectors, including healthcare, education, law and industrial automation. In view of the rapid development of LLMs, it can be assumed that LLMs will soon also be deployed on edge and handheld devices. Several studies have indicated that the current iteration of 5G networks will not be able to support a plethora of services offered by LLMs. Therefore, researchers are working intensively on the next iteration of communication systems, i.e. Sixth generation (6G), to meet the above requirements [2]. Furthermore, as AI is an integral part of 6G systems, it is assumed that LLMs will be used intrinsically to optimize resources and performance while enabling human-centric customized services to users.
@@
Figure 1: Network for LLMs illustration in 6G with examples of smart homes and emergency services. The LLMs can be used by the central cloud and shared with the 6G Edge Cloud. The 6G edge cloud then share either the parameters to the radio access networks or users to fine-tune the network for personalization, or it shres the cached version of the model to provide a specific service.
@@
The standardization process towards 6G systems is already progressing steadily. It is assumed that the evolution of the communication system will support distributed AI both for edge devices and within the mobile network [2]. Although many researchers argue that the edge devices will not support the use of LLMs, but with continuous breakthroughs in the field of AI, support for edge devices can be extended through distributed learning techniques such as federated learning (FL) and split learning (SL) [3]. In addition, quantization and training can be used to fine-tune an LLM on the edge devices. As proposed in [4], an LLM with 65 billion parameters can be fine-tuned with quantized low-rank adapters (QLoRA) on a downstream task within a single day, achieving comparable performance compared to other state-of-the-art (SOTA) LLMs. It can be assumed that the convergence of quantized networks, LLMs and 6G Multi-Access Edge Computing (MEC) could result in many innovative applications. Researchers have already begun to explore the mutual convergence of LLMs and 6G MEC networks, calling them “LLMs for networks” and “networks for LLMs” respectively. We illustrate a network for LLMs that corresponds to the vision of the Network for AI (NetAI) with respect to a 6G communication system in Figure 1. The NetAI vision supports LLM deployments related to the MEC architecture [2]. Various services such as smart homes, healthcare, education, emergencies, mission-critical applications and finance can be supported with the network for LLMs.
@@
Most research today focuses on the integration of LLMs and communication networks, which would undoubtedly bring unprecedented advances and technological innovation. However, one aspect of this technological progress is being overlooked, namely the security aspect. With all the possibilities and potential of LLMs and the 6G ecosystem, we have to ask ourselves, are LLMs trustworthy? Despite their ability to fine-tune to the downstream task, LLMs are deep neural networks that are vulnerable to privacy attacks, such as model inversion, model poisoning and membership leakage [3, 5]. The growing landscape of LLMs and their integration into communication systems therefore makes it necessary to address security concerns and the development of trustable AI encoders to safeguard the integrity of users and services in 6G systems. To the best of our knowledge, the studies have not explored the security vulnerabilities in network for LLMs (Net4LLMs), which subsequently leaves us defenseless against such attacks as we progress towards the Net4AI vision.
@@
Figure 2: Training, Fine-tuning, and customization strategies for LLMs in MEC 6G framework.
@@ II Related Works
Recently, we have observed a plethora of advances in the field of LLMs that would be a revolution in the field of communication networks, especially in the design and development of 6G networks. Some studies have already explored and demonstrated the significance of LLMs for potential 6G applications. For example, the study in [7] proposed NetGPT, which enables personalized services to users through generative networks while handling comprehensive network intelligence and cloud collaboration in real time.Xu et al. [8] focused on the data privacy in 6G communication systems using LLMs. The study proposed to design LLM agents based on the principle of split learning by distributing LLMs for different roles across edge devices to make user interaction efficient and collaborative. Their results show that the split learning setting was effective in improving the communication efficiency while offloading the tasks that are complex in nature to the servers for constructing global LLMs. The study in [9] emphasised that the newer LLMs must offer multimodal services, i.e. they must handle image, text and audio data in order to offer automated services. Therefore, the deployment of LLM agents in the cloud could pose challenges in terms of data privacy, high bandwidth costs and long response times. However, MEC based on 6G communication systems can address the above problems in an effective way. Lin et al. [10]also proposed a split learning framework for the deployment of LLMs in 6G networks. However, their work focused on the efficiency and effectiveness of LLMs in terms of parameter sharing, quantization, and efficient fine-tuning rather than data or model security. Nguyen et al. [5] highlighted the advantages of using LLMs in 6G networks while exploring the security vulnerabilities from an adversarial point of view. However, the discussion of model security was very abstract and brief, focusing more on the attacks on services. The study also suggested the use of blockchain technology to avoid the security threats associated with LLMs and 6G networks. To the best of our knowledge, none of the studies investigated a specific model-based attack scenario related to pre-trained AI encoders and 6G networks.
III Training/Fine-tuning LLMs in 6G
@@
The training strategies for fine-tuning LLMs in 6G networks are shown in Figure 2. The strategies are presented in accordance with the MEC framework. The cloud layer can leverage the pre-trained AI encoders of LLMs for any of the training modes, but the downstream task would mostly be generalized when passed to the edge and user layers. However, at the edge and user layer, the customization of the LLMs can be done based on the context and resources. As can be seen from the training modes, fine-tuning the pre-trained encoder and classifier requires large amounts of computational resources. The computational resources decrease significantly when moving to the edge and user layer. In this regard, the edge and user layers can at most fine-tune/adapt the LLMs based on their application by keeping the pre-trained encoder frozen and using the final layers to update the parameters. Alternatively, they can simply freeze the entire LLM and extract only the output embeddings that are used as representative features for training deep neural networks or shallow learning methods.
@@
III-A Security Issues in LLMs and 6G
Several studies have now highlighted the security concerns related to the behavior, architecture and design of LLM. The security concerns arise from the complexity of LLMs and the challenges associated with their deployment and training process. In addition, backdoor attacks are possible in LLMs that cannot be overcome with conventional security measures. These backdoor attacks are applicable to LLMs that are fine-tuned in a supervised manner and trained with adversarial learning or reinforcement learning. The different types of attacks in LLMs deployed within 6G networks are defined below.
@@ IV-A Threat Model
Before we define the threat model, we make some assumptions. We assume that the LLM is pre- trained on a large dataset capable of transforming the input (text) into embeddings. Using the pre-trained AI encoder, a downstream task is fine-tuned by a customized dataset for a specific application in 6G networks using optimization algorithms and a predefined loss function. The fine-tuned model is then able to transform the input into embeddings or classification probabilities accordingly (which differ from the original, pre-trained AI encoder). We define the attacker’s purpose in this scenario as a dichotomous classification problem, where the goal is to determine whether the input provided to the pre-trained AI encoder is a member or non-member of the training dataset used for the subsequent task. In general, existing studies assume two dimensions of an attacker’s background knowledge when considering membership inference attacks.
The first dimension assumes a black box attack, which means that the attacker has no prior knowledge of the pre-trained AI encoder architecture, but the attacker has access to the model that has been trained for the downstream task. This is considered the most realistic scenario, as in 6G AI is used as a service and the models adapted for the downstream task would be directly available to the public. The second dimension assumes that the attacker has access to a very small subset of the member training data, which can be used to create an auxiliary dataset. The auxiliary dataset can then be used to train the attack model. Studies have shown that such assumptions can be true if one infers the location and makes an educated guess about the service used in a particular area [11]. With the large plethora of diverse data available on the Internet, it is reasonable to assume that the attacker can gather meaningful data to create a shadow model for lodging membership inference attack that corresponds to a real-world environment. Considering the two dimensions, we assume in this study that the attacker has access to the downstream task model and has some knowledge of the application, which is taken into account by the pre-trained AI encoder of the LLM.
@@ V Experimental analysis
The success rate of the attack raises serious concerns about the trustworthiness of LLMs and pre-trained models in 6G networks. We repeat the above experiment with a reduced auxiliary dataset, i.e., we use Yelp Review/AG’s News/SST for training the attack model without considering CoNLL2003 and vice versa to observe the results. The results for this experiment are shown in Figure 4. It can be seen that the performance is still above the random guess. Furthermore, the performance degradation is about 0.12/0.096 for CoNNL2003, about 0.07/0.095 for Yelp Review, about 0.06/0.075 for AG’s News and about 0.105/0.074 for SST, using ALBERT and RoBERTa, respectively. The attack performance still reaches a maximum of 0.83 F1 score if the attacker only has access to the fine-tuned model for the downstream task, but does not use the part of the same datasets. This behavior confirms our assumption that the pre-trained language models memorize the data and behave differently with member and non-member data.
@@ VI Enabling Trust with Pre-trained AI encoders
Membership leakage and membership inference attacks have been extensively studied in the context of computer vision and image modality. Defenses against such attacks therefore include adversary regularization, differential privacy, data augmentation, adding noise to images, intentional attacks, encryption techniques, and others [3, 11]. Some of the above techniques are difficult to perform in textual modality, such as addition of noise and intentional attack initialization. Such actions can also degrade the performance of LLMs in 6G networks. Adding noise to either the data or confidence scores for classification can be used as a defense mechanism. However, studies suggest that such techniques degrade the performance of the downstream task [5].
@@
The second defense is based on the intuition “confidence is defined by trust” (a quote from Patrick Mosher). In this case, however, we would look at confidence and trust from the perspective of AI. We propose to use a trust evaluation module at the edge layer that could evaluate the trust of the pre-trained AI encoder or a fine-tuned LLM with a predefined metric. One of the examples of such a trust evaluation is as follows, assuming the fine-tuned LLM is trained for medical emergency services using a 6G network.
@@
Given the scenario described above, a trust module might favor the first model as it is less vulnerable to membership leakage attack. The idea is simply to prioritize a fine-tuned LLM for 6G services based on trustworthiness and confidence scores.
@@ VII Open Issues, Challenges and Future Directions
The integration of 6G and LLMs can be seen as task-oriented communication services, where integration is achieved by utilizing resources from the communication infrastructure, the edge and mobile devices. In return, users receive LLM agents that can perform certain actions, generate data or call application programming interface (API) functions. As already indicated, such integration can lead to security vulnerabilities, including the theft of personal information and more. We have emphasised the importance of a trust module for considered integration, but designing such a trust module can present some challenges. These challenges include, but are not limited to, the following.
@@
(ii) Multimodal LLMs: In this paper, we have focused on the LLMs that only work with text data. In reality, users opt for multimodal LLMs that can generate text, images and audio. Each modality has its own security issues when it comes to pre-trained AI encoders. However, it would be quite a challenge to design a trust module for 6G and LLMs that is suitable for different data modalities.
(iii) User Privacy: Similar to the design of trust modules for different data modalities, training processes, architectural modifications, and encryption techniques must be used to improve user privacy. Various attacks such as model inversion, model poisoning, gradient leakages, and adversarial attacks can be used by attackers to disrupt the services and steal users’ private information in 6G networks. Therefore, in addition to data modality, the trust module must also defend against various privacy attacks.

### v2

# Pathway to Secure and Trustworthy ZSM for LLMs: Attacks, Defense, and Opportunities

@@ Abstract
Recently, large language models (LLMs) have been gaining a lot of interest due to their adaptability and extensibility in emerging applications, including communication networks. It is anticipated that zero touch network and service management (ZSM) will be able to support mobile edge computing networks and LLMs as a service, as they aim to provide network automation and service management without requiring manual intervention. However, LLMs are vulnerable to data and model privacy issues that affect the trustworthiness of LLMs to be deployed for user-based services. In this paper, we explore the security vulnerabilities associated with fine-tuning LLMs in ZSM in particular the membership inference attack. We define the characteristics of an attack network that can perform a membership inference attack if the attacker has access to the fine-tuned model for the downstream task. We show that the membership inference attacks are effective for any downstream task, which can lead to a personal data breach when using LLM as a service. The experimental results show that the attack success rate of maximum 92% can be achieved on named entity recognition task. Based on the experimental analysis, we discuss possible defense mechanisms and present possible research directions to make the LLMs more trustworthy in the context of 6G networks.
@@ I Introduction
The emergence of attention networks has been a stepping stone for transformer architectures, which also led to the introduction of large language models (LLMs). More recently, LLMs are seen as the most significant advance in the field of artificial intelligence (AI) and a potential pathway to artificial general intelligence (AGI) [1]. Every tech giant is in a race to advance in the field of LLMs by leveraging generative AI (GAI). Notable examples of LLMs from the tech giants are GPT-4 from OpenAI, LLaMA-3 from Meta and PALM from Google. However, there are also new players in this field that surpass the performance of the LLMs mentioned above. These include Mistral (in collaboration with NVIDIA), DCLM from Apple, xLAM from Salesforce, v2 chat from Deepseek, Groq, Claude, SmolLM and many more. These LLMs are trained on diverse and large amounts of datasets scraped or curated from the Internet. Some LLMs focus on increasing model size, such as GPT, while others find new ways to improve the generalization of LLMs through data curation, model quantization, and innovative techniques. Examples of such LLMs are Claude, LLaMA, DCLM and Groq, which have recently outperformed GPT on various language tasks. In continuation of the above-mentioned advances in LLMs, several enterprises are leveraging pre-trained encoders of LLMs to varying degrees to develop their own customized solutions for various applications and sectors, including healthcare, education, law and industrial automation. In view of the rapid development of LLMs, it can be assumed that LLMs will soon also be deployed on edge and handheld devices. Several studies have indicated that the current iteration of networks will not be able to support a plethora of services offered by LLMs. Therefore, researchers are working intensively on the next iteration of communication systems, i.e. Sixth generation (6G) and zero touch network and service management (ZSM), to meet the above requirements [2]. Furthermore, as AI is an integral part of ZSM, it is assumed that LLMs will be used intrinsically to optimize resources and performance while enabling human-centric customized services to users.
@@
Figure 1: Network for LLMs illustration in ZSM with examples of smart homes and emergency services. The LLMs can be used by the central cloud and shared with the 6G Edge Cloud. The 6G edge cloud then share either the parameters to the ZSM network or users to fine-tune the network for personalization, or it shares the cached version of the model to provide a specific service.
@@
The standardization process towards 6G systems and ZSM networks is already progressing steadily. It is assumed that the evolution of the aforementioned systems will support distributed AI both for edge devices and within the mobile network [2]. Although many researchers argue that the edge devices will not support the use of LLMs, but with continuous breakthroughs in the field of AI, support for edge devices can be extended through distributed learning techniques such as federated learning (FL) and split learning (SL) [3]. In addition, quantization and training can be used to fine-tune an LLM on the edge devices. As proposed in [4], an LLM with 65 billion parameters can be fine-tuned with quantized low-rank adapters (QLoRA) on a downstream task within a single day, achieving comparable performance compared to other state-of-the-art (SOTA) LLMs. It can be assumed that the convergence of quantized networks, LLMs, 6G, and ZSM-based Multi-Access Edge Computing (MEC) could result in many innovative applications. Researchers have already begun to explore the mutual convergence of LLMs, ZSM, and MEC networks, calling them “LLMs for networks” and “networks for LLMs” respectively. We illustrate a network for LLMs that corresponds to the vision of the Network for AI (NetAI) with respect to an integrated 6G-ZSM communication system in Figure 1. The NetAI vision supports LLM deployments related to the MEC architecture [2]. Various services such as smart homes, healthcare, education, emergencies, mission-critical applications and finance can be supported with the network for LLMs.
@@
Most research today focuses on the integration of LLMs and communication networks, which would undoubtedly bring unprecedented advances and technological innovation. However, one aspect of this technological progress is being overlooked, namely the security aspect. With all the possibilities and potential of LLMs and the ZSM ecosystem, we have to ask ourselves, are LLMs trustworthy? Despite their ability to fine-tune to the downstream task, LLMs are deep neural networks that are vulnerable to privacy attacks, such as model inversion, model poisoning and membership leakage [3, 5]. The growing landscape of LLMs and their integration into communication systems therefore makes it necessary to address security concerns and the development of trustable AI encoders to safeguard the integrity of users and services in ZSM networks. To the best of our knowledge, the studies have not explored the security vulnerabilities in network for LLMs (Net4LLMs), which subsequently leaves us defenseless against such attacks as we progress towards the Net4AI vision.
@@
Figure 2: Training, Fine-tuning, and customization strategies for LLMs in ZSM-based MEC framework.
@@ II Related Works
Recently, we have observed a plethora of advances in the field of LLMs that would be a revolution in the field of communication networks, especially in the design and development of ZSM networks. Some studies have already explored and demonstrated the significance of LLMs for potential ZSM applications. For example, the study in [7] proposed NetGPT, which enables personalized services to users through generative networks while handling comprehensive network intelligence and cloud collaboration in real time.Xu et al. [8] focused on the data privacy in 6G communication systems using ZSM and LLMs. The study proposed to design LLM agents based on the principle of split learning by distributing LLMs for different roles across edge devices to make user interaction efficient and collaborative. Their results show that the split learning setting was effective in improving the communication efficiency while offloading the tasks that are complex in nature to the servers for constructing global LLMs. The study in [9] emphasized that the newer LLMs must offer multimodal services, i.e. they must handle image, text and audio data in order to offer automated services. Therefore, the deployment of LLM agents in the cloud could pose challenges in terms of data privacy, high bandwidth costs and long response times. However, MEC based on ZSM networks can address the above problems in an effective way. Lin et al. [10] also proposed a split learning framework for the deployment of LLMs in 6G-based ZSM networks. However, their work focused on the efficiency and effectiveness of LLMs in terms of parameter sharing, quantization, and efficient fine-tuning rather than data or model security. Nguyen et al. [5] highlighted the advantages of using LLMs in ZSM networks while exploring the security vulnerabilities from an adversarial point of view. However, the discussion of model security was very abstract and brief, focusing more on the attacks on services. The study also suggested the use of blockchain technology to avoid the security threats associated with LLMs and ZSM networks. To the best of our knowledge, none of the studies investigated a specific model-based attack scenario related to pre-trained AI encoders and ZSM networks.
III Training/Fine-tuning LLMs in ZSM
@@
The training strategies for fine-tuning LLMs in ZSM networks are shown in Figure 2. The strategies are presented in accordance with the MEC framework. The cloud layer can leverage the pre-trained AI encoders of LLMs for any of the training modes, but the downstream task would mostly be generalized when passed to the edge and user layers. However, at the edge and user layer, the customization of the LLMs can be done based on the context and resources. As can be seen from the training modes, fine-tuning the pre-trained encoder and classifier requires large amounts of computational resources. The computational resources decrease significantly when moving to the edge and user layer. In this regard, the edge and user layers can at most fine-tune/adapt the LLMs based on their application by keeping the pre-trained encoder frozen and using the final layers to update the parameters. Alternatively, they can simply freeze the entire LLM and extract only the output embeddings that are used as representative features for training deep neural networks or shallow learning methods.
@@
III-A Security Issues in LLMs and ZSM
Several studies have now highlighted the security concerns related to the behavior, architecture and design of LLM. The security concerns arise from the complexity of LLMs and the challenges associated with their deployment and training process. In addition, backdoor attacks are possible in LLMs that cannot be overcome with conventional security measures. These backdoor attacks are applicable to LLMs that are fine-tuned in a supervised manner and trained with adversarial learning or reinforcement learning. The different types of attacks in LLMs deployed within ZSM networks are defined below.
@@ IV-A Threat Model
Before we define the threat model, we make some assumptions. We assume that the LLM is pre- trained on a large dataset capable of transforming the input (text) into embeddings. Using the pre-trained AI encoder, a downstream task is fine-tuned by a customized dataset for a specific application in ZSM networks using optimization algorithms and a predefined loss function. The fine-tuned model is then able to transform the input into embeddings or classification probabilities accordingly (which differ from the original, pre-trained AI encoder). We define the attacker’s purpose in this scenario as a dichotomous classification problem, where the goal is to determine whether the input provided to the pre-trained AI encoder is a member or non-member of the training dataset used for the subsequent task. In general, existing studies assume two dimensions of an attacker’s background knowledge when considering membership inference attacks.
The first dimension assumes a black box attack, which means that the attacker has no prior knowledge of the pre-trained AI encoder architecture, but the attacker has access to the model that has been trained for the downstream task. This is considered the most realistic scenario, as in ZSM, AI is used as a service and the models adapted for the downstream task would be directly available to the public. The second dimension assumes that the attacker has access to a very small subset of the member training data, which can be used to create an auxiliary dataset. The auxiliary dataset can then be used to train the attack model. Studies have shown that such assumptions can be true if one infers the location and makes an educated guess about the service used in a particular area [11]. With the large plethora of diverse data available on the Internet, it is reasonable to assume that the attacker can gather meaningful data to create a shadow model for lodging membership inference attack that corresponds to a real-world environment. Considering the two dimensions, we assume in this study that the attacker has access to the downstream task model and has some knowledge of the application, which is taken into account by the pre-trained AI encoder of the LLM.
@@ V Experimental analysis
The success rate of the attack raises serious concerns about the trustworthiness of LLMs and pre-trained models in ZSM networks. We repeat the above experiment with a reduced auxiliary dataset, i.e., we use Yelp Review/AG’s News/SST for training the attack model without considering CoNLL2003 and vice versa to observe the results. The results for this experiment are shown in Figure 4. It can be seen that the performance is still above the random guess. Furthermore, the performance degradation is about 0.12/0.096 for CoNNL2003, about 0.07/0.095 for Yelp Review, about 0.06/0.075 for AG’s News and about 0.105/0.074 for SST, using ALBERT and RoBERTa, respectively. The attack performance still reaches a maximum of 0.83 F1 score if the attacker only has access to the fine-tuned model for the downstream task, but does not use the part of the same datasets. This behavior confirms our assumption that the pre-trained language models memorize the data and behave differently with member and non-member data.
@@ VI Enabling Trust with Pre-trained AI encoders
Membership leakage and membership inference attacks have been extensively studied in the context of computer vision and image modality. Defenses against such attacks therefore include adversary regularization, differential privacy, data augmentation, adding noise to images, intentional attacks, encryption techniques, and others [3, 11]. Some of the above techniques are difficult to perform in textual modality, such as addition of noise and intentional attack initialization. Such actions can also degrade the performance of LLMs in ZSM networks. Adding noise to either the data or confidence scores for classification can be used as a defense mechanism. However, studies suggest that such techniques degrade the performance of the downstream task [5].
@@
The second defense is based on the intuition “confidence is defined by trust” (a quote from Patrick Mosher). In this case, however, we would look at confidence and trust from the perspective of AI. We propose to use a trust evaluation module at the edge layer that could evaluate the trust of the pre-trained AI encoder or a fine-tuned LLM with a predefined metric. One of the examples of such a trust evaluation is as follows, assuming the fine-tuned LLM is trained for medical emergency services using a ZSM network.
@@
Given the scenario described above, a trust module might favor the first model as it is less vulnerable to membership leakage attack. The idea is simply to prioritize a fine-tuned LLM for ZSM network based on trustworthiness and confidence scores.
@@ VII Open Issues, Challenges and Future Directions
The integration of ZSM and LLMs can be seen as task-oriented communication services, where integration is achieved by utilizing resources from the communication infrastructure, the edge and mobile devices. In return, users receive LLM agents that can perform certain actions, generate data or call application programming interface (API) functions. As already indicated, such integration can lead to security vulnerabilities, including the theft of personal information and more. We have emphasised the importance of a trust module for considered integration, but designing such a trust module can present some challenges. These challenges include, but are not limited to, the following.
@@
(ii) Multimodal LLMs: In this paper, we have focused on the LLMs that only work with text data. In reality, users opt for multimodal LLMs that can generate text, images and audio. Each modality has its own security issues when it comes to pre-trained AI encoders. However, it would be quite a challenge to design a trust module for ZSM and LLMs that is suitable for different data modalities.
(iii) User Privacy: Similar to the design of trust modules for different data modalities, training processes, architectural modifications, and encryption techniques must be used to improve user privacy. Various attacks such as model inversion, model poisoning, gradient leakages, and adversarial attacks can be used by attackers to disrupt the services and steal users’ private information in ZSM networks. Therefore, in addition to data modality, the trust module must also defend against various privacy attacks.